OpenStudy (anonymous):

find the product: (x-y)(x+5y)

OpenStudy (e.mccormick):

Are you familiar with polynomial multiplication? FOIL?

OpenStudy (anonymous):

Yep, FOIL should the the trick, are you familiar with it?

OpenStudy (anonymous):

F - FIRST O - OUTER I - INNER L - LAST

OpenStudy (anonymous):

2x^2+5xy-xy-5y ?

OpenStudy (e.mccormick):

Close, but you missed something on that 5y and then combined like terms.

OpenStudy (anonymous):

5y^2 ? then t combine it would it be just 5 ?

OpenStudy (anonymous):

or 4xy ?

OpenStudy (e.mccormick):

No, , because it is multiplied it is 5yx, so xy+xy+xy+xy+xy-xy

OpenStudy (e.mccormick):

Ah, you fixed it. 4xy, yes.

OpenStudy (e.mccormick):

\(x^2+4xy-5y^2\)

OpenStudy (anonymous):

so 2x^2+4xy-5y^2

OpenStudy (e.mccormick):

/cheer Yep.

OpenStudy (anonymous):

\[(x-y)(x+5y)\] \[x^2+4 x y-5 y^2\]

OpenStudy (e.mccormick):

Wait... whre did that 2 in front come from? Ummm... Noticed it at the last moment. The x is just squared out fron, no 2. K?

OpenStudy (anonymous):

@e.mccormick yeah @xcaseynoel kind of wrote it wrong; she messed up on the 2 in the front...

OpenStudy (e.mccormick):

First \(\times\) First = \(x^2\), Inner \(\times\) Inner = -xy, Outer \(\times\) Outer = \(5xy\), and Last \(\times\) Last = \(-5y^2\) That is what the whole FOIL thing is about.